OpenStudy (anonymous):

A man has 3 unique coins such that the probability of obtaining a head when the coin is flipped is 1 10 , 2 10 and 3 10 , respectively. If he flips each of the 3 coins once, the probability that at least 1 heads appears is p . What is the value of 1000 p ?

hartnn (hartnn):

do you have some idea on how to start ? how to solve 'atleast 1' type of Questions ?

OpenStudy (anonymous):

if i have any idea, then y i should post the question yaar. really no idea

hartnn (hartnn):

okk...so the rule is, P (atleat 1) = 1 - P (none) now you need to find, P(none) = probability that in 3 flips no head appears. ok till here ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

probability that in 3 flips no head appears = probaility that head doesn't appear (means tail appears) on 1st flip AND probaility that head doesn't appear (means tail appears) on 2nd flip AND probaility that head doesn't appear (means tail appears) on3rd flip right ?

OpenStudy (anonymous):

yup

hartnn (hartnn):

so, if you flip 1st coin, what is the probability of getting tail(not getting head) ?

OpenStudy (anonymous):

no idea

hartnn (hartnn):

is this your 1st probability Question ?

OpenStudy (anonymous):

1/2

hartnn (hartnn):

that woud be true if the coin was unbaised, but here the coin is biased , and we r given P (Head) , so, P (Tail) = 1- P(Head) =... ?

OpenStudy (anonymous):

1-p

hartnn (hartnn):

no..for 1st coin P (Head) = 1/10 given so, P (Tail) = 1-1/10 = 9/10 got this ?

OpenStudy (anonymous):

k..got it

hartnn (hartnn):

for 2nd and 3rd coin, what will be P (Tail) =??

OpenStudy (anonymous):

1-2/10 1-3/10

OpenStudy (anonymous):

8/10 &7/10

hartnn (hartnn):

correct, probability that in 3 flips no head appears = probaility that head doesn't appear (means tail appears) on 1st flip AND probaility that head doesn't appear (means tail appears) on 2nd flip AND probaility that head doesn't appear (means tail appears) on3rd flip =P (Tail) for 1st coin*P (Tail) for 2ndcoin*P (Tail) for 3rd coin = ..... ?

OpenStudy (anonymous):

(9/10)*(8/10)*(7/10)

hartnn (hartnn):

yes, that is p p = (9/10)*(8/10)*(7/10) so, 1000p = 1000 * (9/10)*(8/10)*(7/10) =...?

OpenStudy (anonymous):

9*8*7

hartnn (hartnn):

\(\checkmark \) i hope you understand all steps..

OpenStudy (anonymous):

yup thank u @hartnn

hartnn (hartnn):

welcome ^_^